# Benford behavior resulting from stick and box fragmentation processes

### v2

@@ 2.6. Evaluation of sums within and across intervals
$\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$. (2.21)
By Proposition 2.5 and Section 2.5, we have
@@
$\ =\ (b-a)N^{\delta}\binom{k(N)}{k_{1,\ell}}+O\left(\binom{k(N)}{k_{1,\ell}}N^{\delta\left(1-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right).$ (2.22)
@@
We justify the error term in the last line of (2.6). Since $\kappa$ is the irrationality exponent of an irrational number, then $\kappa>1$ (see [BBS]). Also, $\epsilon^{\prime}>0$, so $1<-1/\kappa+\epsilon^{\prime}$. Moreover, $1/\kappa+\epsilon^{\prime}<0$ by assumption. Combining these with the fact that $\delta\in(0,\epsilon/10)$ gives the dominant error term in the last line of (2.6).
@@
Now, we count the number of stick lengths $A_{k_{1,\ell,i},k_{2},\dots,k_{m}}$ such that $\log_{B}(A_{k_{1,\ell,i},k_{2},\dots,k_{m}})$ mod 1 is in $(a,b)$ over all the intervals. By the truncation in Section 2.3 the main term comes from the sum over $\ell$ from $\sqrt{k(N)}/2$ to $\sqrt{k(N)}/2$. The number is given by
$\sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}\ =\ \sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}+\sum_{|\ell|>\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$. (2.23)
Using (2.14), we can provide an upper bound on the second sum on the RHS of (2.23):
@@
$\sum_{|\ell|>\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$ $\ \leq\ \mathbb{P}\left(\left|k_{1}-\frac{k(N)}{2}\right|\ \geq\ \frac{N^{\delta}\sqrt{k(N)}}{2}\right)\cdot\sum_{0\leq k_{1}\leq N}\binom{k(N)}{k_{1}}$
$\ =\ \mathbb{P}\left(\left|k_{1}-\frac{k(N)}{2}\right|\ \geq\ \frac{N^{\delta}\sqrt{k(N)}}{2}\right)\cdot 2^{k(N)}\ =\ O\left(\frac{2^{k(N)}}{N^{2\delta}}\right).$ (2.24)
On the other hand, we apply (2.6) to the first sum on the RHS of (2.23):
@@
$\sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}\ =\ \sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}(b-a)N^{\delta}\binom{k(N)}{k_{1,\ell}}+O\left(\sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\binom{k(N)}{k_{1,\ell}}N^{\delta\left(1-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right).$ (2.25)
By (2.6) and (5.17) and (5.18) of [B–], by choosing $q=N^{\delta}$, we have
@@
$\sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$ $\ =\ (b-a)2^{k(N)}+O\left(2^{k(N)}N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right).$ (2.26)
Substituting (2.6) and (2.26) into (2.23) yields
@@
$\sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}\ =\ (b-a)2^{k(N)}+O\left(2^{k(N)}N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right).$ (2.27)
@@
Finally, we sum over all $k(N),k_{3},k_{4},\dots,k_{m}$, which gives us the total number of lengths $A_{k_{1},k_{2},\dots,k_{m}}$ such that $\log_{B}(A_{k_{1},k_{2},\dots,k_{m}})$ mod 1 is in $(a,b)$. We break the sum based on whether $k(N)\geq N^{\epsilon}$ or not. For $k(N)\geq N^{\epsilon}$, we apply (2.27):
@@
$\quad+O\left(N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\sum_{\begin{subarray}{c}0\leq k(N),k_{3},k_{4},\dots,k_{m}\leq N\\ k(N)+k_{3}+k_{4}+\cdots+k_{m}=N\\ k(N)\geq N^{\epsilon}\end{subarray}}2^{k(N)}\binom{N}{k(N),k_{3},k_{4},\dots,k_{m}}\right)$. (2.28)
Note that by Lemma 2.3
@@
$\ =\ m^{N}+O((m-2)^{N}N^{N^{\epsilon}+2}),$ (2.29)
where in the last line, we apply Proposition 2.4. Substituting (2.6) into (2.6) gives
@@
$\ =\ (b-a)m^{N}+O\left((m-2)^{N}N^{N^{\epsilon}+2}+m^{N}N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}+(m-2)^{N}N^{N^{\epsilon}+2}N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right)$
$\ =\ (b-a)m^{N}+O\left(m^{N}N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right).$ (2.30)
For $k(N)<N^{\epsilon}$, by Lemma 2.7 and Proposition 2.4 we have
@@
$\ =\ \sum_{\begin{subarray}{c}0\leq k_{1},k_{2},\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}\ =\ O\left((m-2)^{N}N^{N^{\epsilon}+2}\right).$ (2.31)
@@
$\ =\ (b-a)m^{N}+O\left(N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}m^{N}+N^{N^{\epsilon}+2}(m-2)^{N}\right)$
$\ =\ (b-a)m^{N}+O\left(N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}m^{N}\right),$ (2.32)
@@
$F_{N}(a,b)\ =\ b-a+O\left(N^{\delta\left(-\frac{1}{\kappa}+\epsilon^{\prime}\right)}\right).$ (2.33)
We have shown that if $\log(p_{1}/p_{2})$ has a finite irrationality exponent $\kappa$, then the logarithm of the stick lengths is equidistributed mod 1, and thus the stick lengths converge to the Benford distribution. If $\log(p_{1}/p_{2})$ has infinite irrationality exponent, the calculation is exactly the same, except that we start with the error term $o(N^{\delta})$ instead of $O(N^{\delta\left(1-\frac{1}{\kappa}+\epsilon^{\prime}\right)})$ in Section 2.5, and so we end up with the $F_{N}(a,b)\ =\ b-a+o(1)$. Thus, this establishes the strong Benford behavior of fixed multi-proportion 1-dimensional stick fragmentation process when $y_{i}\not\in\mathbb{Q}$ for some $1\leq i\leq m-1$. ∎
3. Proof of Theorem 1.13
3.1. Preliminary
In this section, we prove Theorem 1.13, i.e., resolve Conjecture 4.1 of [BDMMM] that we mentioned in Section 1 under some mild conditions. Let us first recall the conjecture.
Conjecture 3.1.
@@ 3.1. Preliminary
[BDMMM] has verified the conjecture in the case of $d=1$ and arbitrary $m$ for i.i.d. log-uniform distributions, i.e., the proportion at which each dimension is cut is $P_{i}$, where $\log P_{i}\sim\textup{Uniform}(a,b)$ and $a<b\leq 0$. They decide to normalize $P_{i}$ to $\textup{Uniform}(-\sqrt{3},\sqrt{3})$ with mean 0 and variance 1, which simplifies the calculation later on and does not affect the nature or the solution to the problem. After stage $N$, the side length of the $i$th dimension of the box, after taking logarithm and normalizing, is
$Z_{i}^{(N)}\ :=\ \frac{\log_{B}(P_{i}^{(1)}\cdots P_{i}^{(N)})-N\mu_{P}}{\sqrt{N}\sigma_{P}}\ =\ \frac{\log_{B}(P_{i}^{(1)}\cdots P_{i}^{(N)})}{\sqrt{N}}.$ (3.1)
By the CLT, $Z_{i}^{(N)}$ converges to $N(0,1)$. Moreover, Lemma 3.1 states that if $f_{Z_{i}^{(N)}}(z)$ is the probability density function of $Z_{i}^{(N)}$ and $\varphi(z)$ is the probability density function of $N(0,1)$, then for all $z\in[-\sqrt{3}N,\sqrt{3}N]$ and any $\epsilon>0$, we have
$f_{Z_{i}^{(N)}}(z_{i})\ =\ \varphi(z_{i})+O(N^{-1+4\epsilon}),$ (3.2)
where
$\varphi(x)\ :=\ \frac{1}{\sqrt{2\pi}}e^{-x^{2}/2}.$ (3.3)
Due to the presence of the error term, to establish the strong Benford behavior of the process, they need to rely on the fact that the distributions for the cuts are finitely supported with uniformly bounded supports. They also need to rely on the fact that the exponent of the error term is less than 1/2-1/2 for sufficiently small $\epsilon>0$, which allows the error term from integrating $f_{Z^{(N)}_{i}}(z)$ over the support to be asymptotically 0.
For the rest of this paper, we prove Conjecture 4.1 of [BDMMM] for any $m\geq 2$ and $d\leq m$. Physically, this means that we consider any arbitrary dimensional volume of faces of that dimension of the $m$-dimensional box. We assume that the $\log(P_{i})$ are i.i.d., have mean 0 and variance 1, are finitely supported with support [C,C][-C,C], and that the probability density function (PDF) of $Z_{i}^{(N)}$ is $f_{Z_{i}^{(N)}}(z_{i})=\varphi(z_{i})+A(z_{i})$, where $A(z_{i})=O(N^{-1/2-\delta})$ for some $\delta>0$. We also assume that the cumulative density function (CDF) of $Z_{i}^{(N)}$ is given by $F_{Z_{i}^{(N)}}(z_{i})=\Phi(z_{i})+B(z_{i})$, where $\Phi$ is the CDF of $N(0,1)$ and $B(z_{i})=O(N^{-\delta})$. We prove that the maximum $d$-dimensional volume $\mathfrak{m}_{d}^{(N)}$ of a $d$-dimensional face of the $m$-dimensional box converges to strong Benford behavior. Once we prove this statement, then by the maximum criterion established in [BDMMM], we have that the $d$-dimensional volume $V_{d}^{(N)}$ of the $d$-dimensional faces of the $m$-dimensional box also converges to strong Benford behavior.
The key idea behind our proof is the use of the Mean Value Theorem and a recursive method of change of variables and consistent cancellation that provide upper bounds on the error terms in terms of Gaussian densities. We shall explore this more in Section 3.6.
To begin, we use knowledge from order statistics to find the PDF of $\mathfrak{m}_{d}^{(N)}$.
Definition 3.2.
@@
Let $S_{i}^{(N)}$ be the side length of the $i\textsuperscript{th}$ dimension of the $m$-dimensional box after stage $N$. Then the maximum volume of a $d$-dimensional face is the product of the longest $d$ sides of the box, i.e., $\mathfrak{m}_{d}^{(N)}=\prod_{i=1}^{d}S_{(m+1-i)}^{(N)}$. We define
$Y^{(N)}\ :=\ \frac{\log_{B}(\mathfrak{m}_{d}^{(N)})}{\sqrt{N}}\ =\ \sum_{i=1}^{d}\frac{\log_{B}(S^{(N)}_{(m+1-i)})}{\sqrt{N}}\ =\ \sum_{i=1}^{d}Z_{(m+1-i)}^{(N)},$ (3.5)

### v3

@@ 2.6 Evaluation of sums within and across intervals
$\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$. (33)
By Proposition 3 and Section 2.5, we have
@@
$\ =\ (b-a)N^{\delta}\binom{k(N)}{k_{1,\ell}}+O\left(\binom{k(N)}{k_{1,\ell}}N^{\delta\left(1-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\right).$ (34)
@@
We justify the error term in the last line of (2.6). Since $\kappa$ is the irrationality exponent of an irrational number, then $\kappa\geq 2$ (see [2]). Also, $\epsilon^{\prime}>0$, so $1<-1/(\kappa-1)+\epsilon^{\prime}$. Moreover, for some proper choice of $\epsilon^{\prime}$, we have $1/(\kappa-1)+\epsilon^{\prime}<0$. Combining these with the fact that $\delta\in(0,\epsilon/10)$ gives the dominant error term in the last line of (2.6).
@@
Now, we count the number of stick lengths $A^{(N)}_{k_{1,\ell,i},k_{2},\dots,k_{m}}$ such that $\log_{B}(A^{(N)}_{k_{1,\ell,i},k_{2},\dots,k_{m}})$ mod 1 is in $(a,b)$ over all the intervals. By the truncation in Section 2.3 the main term comes from the sum over $\ell$ from $\sqrt{k(N)}/2$ to $\sqrt{k(N)}/2$. The number is given by
$\sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}\ =\ \sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}+\sum_{|\ell|>\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$. (35)
Using (24), we can provide an upper bound on the second sum on the RHS of (35):
@@
$\sum_{|\ell|>\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$ $\ \leq\ \mathbb{P}\left(\left|k_{1}-\frac{k(N)}{2}\right|\ \geq\ \frac{N^{\delta}\sqrt{k(N)}}{2}\right)\cdot\sum_{0\leq k_{1}\leq k(N)}\binom{k(N)}{k_{1}}$
$\ =\ \mathbb{P}\left(\left|k_{1}-\frac{k(N)}{2}\right|\ \geq\ \frac{N^{\delta}\sqrt{k(N)}}{2}\right)\cdot 2^{k(N)}\ =\ O\left(\frac{2^{k(N)}}{N^{2\delta}}\right).$ (36)
On the other hand, we apply (2.6) to the first sum on the RHS of (35):
$\sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$ $\ =\ \sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}(b-a)N^{\delta}\binom{k(N)}{k_{1,\ell}}$
@@
$\quad\quad+O\left(\sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\binom{k(N)}{k_{1,\ell}}N^{\delta\left(1-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\right).$ (37)
We know that
$\sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\binom{k(N)}{k_{1,\ell}}$ $\ =\ \sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\binom{k(N)}{k_{1,\ell}}+O\left(\sum_{|\ell|>\sqrt{k(N)}/2}\binom{k(N)}{k_{1,\ell}}\right)$
$\ =\ \sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\binom{k(N)}{k_{1,\ell}}+O\left(\frac{2^{k(N)}}{N^{2\delta}}\right).$ (38)
By (5.17) of [1], by choosing $q=N^{\delta}$, we have
$\sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\binom{k(N)}{k_{1,\ell}}\ =\ \frac{2^{k(N)}}{N^{\delta}}+O\left(2^{k(N)}e^{-3k(N)^{1-2\delta}}\right)$ (39)
Substitute (39) into (2.6) and then substitute the resulting formula into (2.6), we have
@@
$\sum_{\ell=-\sqrt{k(N)}/2}^{\sqrt{k(N)}/2}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}$ $\ =\ (b-a)2^{k(N)}+O\left(2^{k(N)}N^{-\delta}+2^{k(N)}N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\right)$
$\ =\ (b-a)2^{k(N)}+O\left(2^{k(N)}N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\right),$ (40)
where the last line follows from $1<-1/(\kappa-1)+\epsilon^{\prime}<0$. Substituting (2.6) and (2.6) into (35) yields
@@
$\sum_{\ell=-k(N)/(2N^{\delta})}^{k(N)/(2N^{\delta})}\sum_{i\in J_{\ell}(a,b)}\binom{k(N)}{k_{1,\ell,i}}\ =\ (b-a)2^{k(N)}+O\left(2^{k(N)}N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\right).$ (41)
@@
Finally, we sum over all $k(N),k_{3},k_{4},\dots,k_{m}$, which gives us the total number of lengths $A^{(N)}_{k_{1},k_{2},\dots,k_{m}}$ such that $\log_{B}(A^{(N)}_{k_{1},k_{2},\dots,k_{m}})$ mod 1 is in $(a,b)$. We break the sum based on whether $k(N)\geq N^{\epsilon}$ or not. For $k(N)\geq N^{\epsilon}$, we apply (41):
@@
$\quad\quad+O\left(N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\sum_{\begin{subarray}{c}0\leq k(N),k_{3},k_{4},\dots,k_{m}\leq N\\ k(N)+k_{3}+k_{4}+\cdots+k_{m}=N\\ k(N)\geq N^{\epsilon}\end{subarray}}2^{k(N)}\binom{N}{k(N),k_{3},k_{4},\dots,k_{m}}\right)$. (42)
Note that by Lemma 2
@@
$\ =\ m^{N}+O((m-2)^{N}N^{N^{\epsilon}+2}),$ (43)
where in the last line, we apply Proposition 2. Substituting (2.6) into (2.6) gives
@@
$\ =\ (b-a)m^{N}+O\left((m-2)^{N}N^{N^{\epsilon}+2}+m^{N}N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}+(m-2)^{N}N^{N^{\epsilon}+2}N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\right)$
$\ =\ (b-a)m^{N}+O\left(m^{N}N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\right).$ (44)
For $k(N)<N^{\epsilon}$, by Lemma 1 and Proposition 2 we have
@@
$\ =\ \sum_{\begin{subarray}{c}0\leq k_{1},k_{2},\dots,k_{m}\leq N\\ k_{1}+k_{2}+\cdots+k_{m}=N\\ k_{1}+k_{2}<N^{\epsilon}\end{subarray}}\binom{N}{k_{1},k_{2},\dots,k_{m}}\ =\ O\left((m-2)^{N}N^{N^{\epsilon}+2}\right).$ (45)
@@
$\ =\ (b-a)m^{N}+O\left(N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}m^{N}+N^{N^{\epsilon}+2}(m-2)^{N}\right)$
$\ =\ (b-a)m^{N}+O\left(N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}m^{N}\right),$ (46)
@@
$F_{N}(a,b)\ =\ b-a+O\left(N^{\delta\left(-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)}\right).$ (47)
Thus, we have shown that if $\log_{B}(p_{1}/p_{2})$ has a finite irrationality exponent $\kappa$, then the logarithm of the stick lengths is equidistributed mod 1, and thus the stick lengths converge to the Benford distribution. Moreover, the discrepancy between the distribution of the logarithm of the stick length and uniform distribution mod 1 is quantified in terms of the irrationality exponent of $\log_{B}(p_{1}/p_{2})$. Equation (47) tells us that the smaller the irrationality exponent, the smaller the discrepancy is (thanks to the guarantee that the error term is optimal), and the closer the distribution of the stick length is to the Benford distribution. If $\log(p_{1}/p_{2})$ has infinite irrationality exponent, the calculation is exactly the same, except that we start with the error term $o(N^{\delta})$ instead of $O(N^{\delta\left(1-\frac{1}{\kappa-1}+\epsilon^{\prime}\right)})$ in Section 2.5, and so we end up with the $F_{N}(a,b)\ =\ b-a+o(1)$.
Thus, this establishes the strong Benford behavior of fixed multi-proportion 1-dimensional stick fragmentation process when $y_{i}\not\in\mathbb{Q}$ for some $1\leq i\leq m-1$. On a final note, recall from Section 2.1 our choices of the order of $p_{1},p_{2},\dots,p_{m}$ and thus factorization are arbitrary. Hence, to obtain an optimal error term, we simply need to choose an order of $p_{1},p_{2},\dots,p_{m}$ such that $\log_{B}(p_{1}/p_{2})$ is irrational with the least irrationality exponent. Let $\kappa_{0}$ be such irrationality exponent. Then we have
$F_{N}(a,b)\ =\ b-a+O\left(N^{\delta\left(-\frac{1}{\kappa_{0}-1}+\epsilon^{\prime}\right)}\right).$ (48)
3 Proof of Theorem 5
3.1 Preliminary
In this section, we prove Theorem 5, i.e., Conjecture 4.1 of [6], under some conditions. Let us first recall the conjecture.
Conjecture 1.
@@ 3.1 Preliminary
Betti et al. [6] were able to prove the Conjecture for any arbitrary $m\geq 1$ and $d=1$. That is, they showed that under any linear fragmentation process, the length of the longest side of a fragmented box of any arbitrary dimension converges to strong Benford behavior. For the rest of this paper, we prove Theorem 5 for any $m\geq 1$ and $d\leq m$, i.e., we consider the volume of the largest $d$-dimensional face of a fragmented box of any arbitrary dimension $m\geq d$ converges to strong Benford behavior. Our proof relies on the following conditions on the distribution of $P_{i}^{(j)}$:
  1. 1.
    Condition 1: $P_{i}^{(j)}$’s are i.i.d. for all $1\leq i\leq m$ and $1\leq j\leq N$. That is, $P_{i}^{(j)}$’s are all independent and there exists a random variable $P$ such that $P_{i}^{(j)}\sim P$ for all $i,j$;
  2. 2.
    Condition 2: Let the mean and the variance of $\log_{B}(P)$ be $\mu_{P}$ and $\sigma_{P}$. Then $\log_{B}(P)$ has finite centered moments, i.e., $\mathbb{E}[(\log_{B}(P)-\mu_{P})^{k}]<\infty$ for all $k\geq 1$ and the characteristic function of $(P-\mu_{P})/\sigma_{P}$ admits Taylor expansion. Moreover, the variance is nonzero;
  3. 3.
    Condition 3: Let $f(x)$ be the probability density function of $\log_{B}(P)$, then we have $\textup{ess sup}_{x}f(x)<\infty$;
  4. 4.
    Condition 4: $f(x)$ is differentiable with $\textup{sup}_{x}|f^{\prime}(x)|<\infty$.
These conditions ensure that we are working with “nice” proportion cuts and are satisfied by a wide range of random variables. Moreover, we believe that Theorem 5 should hold in a far more general setting, particular when the $P_{i}^{(j)}$’s have independent but non-identical distributions. For example, as long as $\{P_{i}^{(j)}\}_{j=1}^{N}$ satisfies the conditions for CLT for every $i$, then the distribution of the normalized side lengths of the box are still approximately Gaussian, and the proof techniques developed in this paper should extend without major difficulty. The only technical caveat is perhaps that we would need to use [21] instead of (53) to obtain a formula for the volume of the largest lower dimensional face of the box, since the side lengths are no longer i.i.d.. The formula in [21] is rather complicated–it involves the permanent of a matrix whose entries are simple expressions in terms of the PDF and the CDF of random variables. However, after permanent expansion, it is essentially a linear sum with terms resembling (53). Hence, we believe that our analysis should extend to the general setting where $P_{i}^{(j)}$ are independent but non-identical without too much trouble, but for the clarity of exposition, we decide to restrict ourselves to the i.i.d. setting.
Under these conditions, our goal is then to prove that the volume $\mathfrak{m}_{d}^{(N)}$ of the largest $d$-dimensional face of a $m$-dimensional fragmented box converges to strong Benford behavior. Once we prove this, then by the Maximum Criterion, the volume $V_{d}^{(N)}$ of the $d$-dimensional faces of a $m$-dimensional fragmented box also converges to strong Benford behavior. To begin, let us first introduce the notion of order statistics.
Definition 11.
@@
Let $S_{1}^{(N)},\cdots,S_{m}^{(N)}$ be the side lengths of the $m$-dimensional box after stage $N$. Then the volume of the largest $d$-dimensional face is the product of the longest $d$ side lengths of the box, i.e., $\mathfrak{m}_{d}^{(N)}=\prod_{i=1}^{d}S_{(m+1-i)}^{(N)}$. Since $S_{i}^{(N)}=P_{i}^{(1)}\cdots P_{i}^{(N)}$ for all $i$ and the $P_{i}^{(j)}$’s are between 0 and $1$, then $\mathfrak{m}_{d}^{(N)}$ is asymptotically very small. To make this quantity more convenient to work with, we normalize it as follows. By Condition 1, $P_{i}^{(j)}$’s are i.i.d., which means that $S_{i}^{(N)}$’s are i.i.d. as well. Since $\mathbb{E}[\log_{B}(P)]=\mu_{P}$ and $\textup{Var}[\log_{B}(P)]=\sigma_{P}^{2}$, then for all $i$,
$\mathbb{E}[\log_{B}(S_{i}^{(N)})]\ =\ N\mu_{P}\quad\textup{and}\quad\textup{Var}[\log_{B}(S_{i}^{(N)})]\ =\ N\sigma_{P}^{2}.$ (50)
Now consider
$Z_{i}^{(N)}\ :=\ \frac{\log_{B}(S_{i}^{(N)})-N\mu_{P}}{\sqrt{N}\sigma_{P}}.$ (51)
It is a well-known fact from order statistics that if the order statistics of $(X_{1},\dots,X_{k})$ are $(X_{(1)},\dots,X_{(k)})$, then the order statistics of $(g(X_{1}),\dots,g(X_{k}))$ are $(g(X_{(1)}),\dots,g(X_{(k)}))$ for any monotonically increasing function $g$. Since the order statistics of $(S_{1}^{(N)},\dots,S_{m}^{(N)})$ are $(S_{(1)}^{(N)},\dots,S_{(m)}^{(N)})$ and $g(x):=(\log_{B}(x)-N\mu_{P})/\sqrt{N}\sigma_{P}$ is monotonically increasing, then it follows that the order statistics of $(g(S_{1}^{(N)}),\dots,g(S_{m}^{(N)}))$ are $(g(S_{(1)}^{(N)}),\dots,g(S_{(m)}^{(N)}))$, which are just $(Z^{(N)}_{(1)},\dots,Z^{(N)}_{(m)})$. We are now ready to define the normalized form of $\mathfrak{m}_{d}^{(N)}$ with which we shall use to establish strong Benford behavior:
$Y^{(N)}\ :=\ \frac{\log_{B}(\mathfrak{m}_{d}^{(N)})-d\cdot N\mu_{P}}{\sqrt{N}\sigma_{P}}\ =\ \sum_{i=1}^{d}\frac{\log_{B}(S^{(N)}_{(m+1-i)})-N\mu_{P}}{\sqrt{N}\sigma_{P}}\ =\ \sum_{i=1}^{d}Z_{(m+1-i)}^{(N)}.$ (52)